OpenStudy (darkbluechocobo):

100N is applied over 10 m on a mass of 5kg. The velocity is now 20 m/s. What was the initial velocity.

OpenStudy (anonymous):

F= 100N d = 10 m m= 5 kg Vf= 20 m/s Vi =? F= ma using this formula solve for the acceleration Then, apply one of the 5 kinematics equation: \(\sf V_f^2=V_i^2+2ad\) to solve for \(\sf V_i\) btw, this is the math section (:

OpenStudy (darkbluechocobo):

100/5=a so acceleration=20 and yes sorry s: i posted wrong section

OpenStudy (darkbluechocobo):

20^2=V^2i+2(20)(10)

OpenStudy (anonymous):

don't worry it's okay (: and yes, acceleration value is right yes evaluate that to get \(\sf V_i\)

OpenStudy (darkbluechocobo):

40/400=V^2i

OpenStudy (darkbluechocobo):

400/400*

OpenStudy (anonymous):

\(\sf 20^2=V^2i+2(20)(10) \\ 20^2-2(20)(10) =V_i^2\\V_i= \sqrt{20^2-2(20)(10)}\)

OpenStudy (darkbluechocobo):

so you would subtract 2(20)(10) on both sides to make it negative then sqrt to get just Vi?

OpenStudy (darkbluechocobo):

and the velocity would be 0m/s

OpenStudy (anonymous):

yes, don't divide because the operation between \(\sf V_i^2\) and \(\sf 2ad\) is addition not ,multiplication

OpenStudy (anonymous):

yes :) that would be the initial velocity

OpenStudy (darkbluechocobo):

Thank you :D